RhondaSommer (rhondasommer):

A ladder is leaning up against a 10-foot wall at an angle of elevation of 65°. How long is the ladder? Round your answer to the nearest tenth of a foot. 3.6 feet 4.7 feet 9.1 feet 11.0 feet

OpenStudy (candy13106):

u have to find the hypothese using the Pythagorean theorm and the sin, cosine, or tangent

OpenStudy (candy13106):

do u know how?

RhondaSommer (rhondasommer):

so... cos(-1)(10/x) ... ?

RhondaSommer (rhondasommer):

cos^-1

OpenStudy (candy13106):

u dont have to put a -1

OpenStudy (candy13106):

no positive

RhondaSommer (rhondasommer):

... cos36=10/x

OpenStudy (candy13106):

u know i dont know what u r doing wrong so im going to work this out then help u so can u plz wait

RhondaSommer (rhondasommer):

geez. its not like im being rude im just trying to understand

OpenStudy (candy13106):

sorry i was just being rude

RhondaSommer (rhondasommer):

sin(36)=(x/10)?

OpenStudy (candy13106):

nevermind the first diagram was right sorry about that

RhondaSommer (rhondasommer):

what?

RhondaSommer (rhondasommer):

@Hero help please

OpenStudy (candy13106):

find the tangent not cosine

OpenStudy (candy13106):

tangent is opposite/ adjacent

OpenStudy (candy13106):

tan 65=x/10

OpenStudy (candy13106):

10tan65=x and that will give u the answer got it?

563blackghost (563blackghost):



563blackghost (563blackghost):

this is the visual of the problem....

563blackghost (563blackghost):

Now we know the angle of evelation which is 65 and we know of the opposite side and want to know the hypotenuse which is the length of the ladder so we would use sine... \(\LARGE\color{green}{\sin(65^{o})=\frac{ 10 }{ x }}\) Now to simplify we would divide 10 and sin(65).... \(\LARGE\color{purple}{x=\sin(65^{o}) \div 10}\)

563blackghost (563blackghost):

barely realized this question was answered 5 hours ago.....

563blackghost (563blackghost):

:P

563blackghost (563blackghost):

actually when inputting this problem in your calculator you would first do 10 then divide then sin(65

RhondaSommer (rhondasommer):

so 10/sin(65)

563blackghost (563blackghost):

correct ^^

RhondaSommer (rhondasommer):

okay...

RhondaSommer (rhondasommer):

12.0944038929

563blackghost (563blackghost):

hmmm is your calculator set to degree cause thats the same answer when its set to radian mode.....

RhondaSommer (rhondasommer):

hmmm

RhondaSommer (rhondasommer):

11?

RhondaSommer (rhondasommer):

ish

563blackghost (563blackghost):

yup ^^

563blackghost (563blackghost):

well 11.0337 an on...

563blackghost (563blackghost):

and*

RhondaSommer (rhondasommer):

can you help with more?

563blackghost (563blackghost):

sure ^^